# Budget-Adaptive Adapter Tuning in Orthogonal Subspaces for Continual Learning in LLMs

### v1

@@ Abstract
Large language models (LLMs) often suffer from catastrophic forgetting in continual learning (CL) scenarios, where performance on previously learned tasks degrades severely while training on sequentially arriving tasks. Although pioneering CL approaches using orthogonal subspaces can mitigate task interference, they typically employ fixed budget allocation, neglecting the varying complexity across tasks and layers. Besides, recent budget-adaptive tuning methods for LLMs often adopt multi-stage paradigms that decouple optimization and budget allocation. Such decoupling results in potential misalignment, which hinders those approaches’ practical application in CL scenarios. To address these limitations, we propose OA-Adapter, a novel parameter-efficient approach for continual learning in LLMs that unifies dynamic budget adaptation with orthogonal subspace learning in a single end-to-end training stage. Specifically, OA-Adapter introduces a dynamic bottleneck dimension adaptation mechanism that simultaneously allocates an efficient parameter budget and optimizes task objectives without misalignment. To effectively preserve previously acquired knowledge while coordinating with the dynamic budget allocation, orthogonal constraints are applied specifically between the parameter subspace of the current task and the dynamically allocated parameter subspaces of historical tasks. Experimental results on continual learning benchmarks demonstrate that OA-Adapter outperforms state-of-the-art methods in both accuracy and parameter efficiency, achieving higher average accuracy while using $58.5\%$ fewer parameters on the standard CL benchmark.
@@ 1 Introduction
Recent advances in large language models (LLMs) have transformed artificial intelligence by demonstrating remarkable capabilities across diverse domains, from text generation to logical reasoning [1, 2]. However, real-world deployment demands that LLMs continually adapt to evolving user needs and emerging tasks while retaining previously acquired knowledge—a prerequisite for sustainable lifelong learning [3, 4].
@@
Parameter-efficient fine-tuning (PEFT) methods, such as adapter modules [5] and low-rank adaptation (LoRA) [6], enable task-specific adaptation by updating only $0.01\%-4\%$ of model weights [7]). While originally designed to reduce computational costs for single-task tuning [8, 9], these methods struggle in continual learning with sequentially arriving tasks. Sequentially tuning to arriving tasks induces catastrophic forgetting-severe degradation of performance on prior tasks [10]. One intuitive solution is to store task-specific adapters for each new task. However, this approach consumes substantial storage resources and leads to considerable inflexibility in multi-task deployments. Alternatively, retraining models with archived historical and additional new data necessitates frequent model updates and large data repositories [11]. Both strategies are prohibitively costly and impractical in resource-constrained environments.
@@
To efficiently adapt LLMs to downstream tasks while preserving previously acquired knowledge, researchers proposed continual learning (CL) methods, as discussed in Appendix A.1. However, most CL approaches operate within shared parameter spaces across tasks, which inherently induces cross-task interference [12, 11, 13, 14, 15, 16, 17, 18, 19]. Furthermore, unlike conventional CL benchmarks that typically handle tasks with limited distribution shifts (e.g., incremental image classification), continual learning for LLMs often deals with substantially divergent task distributions, thus significantly amplifying interference when tuning in shared parameter spaces. Some methods further dynamically construct task-specific parameters for new knowledge integration to resolve this problem, but they heavily rely on explicit task identifiers during inference [20, 21, 22, 23, 24].
Recent research efforts in orthogonal subspace learning [25, 26, 27] offer a promising alternative by restricting task-specific updates to mutually orthogonal parameter subspaces, decoupling optimization directions across tasks and thereby eliminating interference and task ID dependency. However, existing methods typically rely on a fixed budget allocation, assigning the same subspace dimensionality to every task and layer. This rigid strategy overlooks the heterogeneity of task complexity and layer-specific adaptation needs, leading to inefficient parameter utilization, allocating excessive resource to simple tasks while under-allocating resources to more complex ones. Such inflexible allocation hinder LLMs’ continuous adaptation capabilities in practice.
@@
To achieve dynamic budget allocation, emerging budget-adaptive PEFT methods like AdaLoRA [28], ElaLoRA [29], and DiffoRA [30] proposed multi-stage paradigms with sequential optimization and budget adjustment phases, as detailed in Appendix A.1. Such decoupled optimization may create misalignment between fine-tuning objectives and budget allocation. Moreover, the inherent complexity of multi-stage designs introduces substantial computational overhead and engineering challenges, limiting their practicality for continual learning systems.
@@
To address these issues, we propose OA-Adapter, a novel parameter-efficient approach for continual learning in LLMs. Instead of manually assigning a fixed budget, OA-Adapter automatically adjusts the parameter budget for each task and layer based on the task difficulty and model capacity. To the best of our knowledge, this is the first work to integrate budget adaptation into parameter-efficient fine-tuning for continual learning in LLMs.
@@
  • We propose OA-Adapter, a novel parameter-efficient approach for continual learning in LLMs that unifies dynamic budget adaptation with orthogonal subspace learning in a single end-to-end training stage.
@@
  • We establish orthogonal constraints between the parameter subspace of the current task and the dynamically allocated parameter subspaces of historical tasks, effectively preserving previously acquired knowledge while coordinating with the dynamic dimension adaptation.
  • Experimental results demonstrate that OA-Adapter outperforms state-of-the-art methods in both accuracy and parameter efficiency. OA-Adapter achieves higher average accuracy with $58.5\%$ fewer parameters on the standard CL benchmark and maintains its advantages on a larger benchmark comprising 15 tasks.
@@ 2 Methodology
In this section, we introduce OA-Adapter, a novel framework for continual learning in LLMs that simultaneously improves parameter efficiency and mitigates catastrophic forgetting in a single end-to-end training stage, as illustrated in Fig. 1. We first describe its architectural design, including core components and computation flow. Then, we analyze the mathematical foundations of the dynamic bottleneck dimension adaptation, demonstrating how trainable thresholds enable bidirectional activation and deactivation of dimensions during a single training phase. Finally, we formalize the orthogonal parameter subspace constraints mechanism and explain how it works in concert with the dynamic bottleneck dimension adaptation to achieve parameter-efficient continual learning.
Figure 1: The OA-Adapter framework for LLM continual learning. Each task-specific OA-Adapter module (task $t$) comprises three core components: (1) a down-projection layer $\mathcal{W}_{1}^{(t)}$, (2) a trainable diagonal mask $\Gamma^{(t)}$ with trainable threshold $\tau^{(t)}$, and (3) an up-projection layer $\mathcal{W}_{2}^{(t)}$. The dynamic masking mechanism enables bidirectional dimension adaptation through activation/deactivation of latent dimensions. Orthogonal subspace constraints are enforced between the column space of the $t$-th task parameters $\mathrm{Col}(\mathcal{W}_{2}^{(t)})$ and the dynamically allocated parameter subspaces of historical tasks $\mathrm{Col}(\widetilde{\mathcal{W}}_{2}^{(s)})$ (for $\ s<t$). Here, $\widetilde{\mathcal{W}}_{2}^{(s)}$ incorporates only the activated dimensions from the $s$-th task.
@@ 2.1 Module Structure
When adapting pre-trained language models (PLMs) to downstream tasks, traditional full-parameter fine-tuning proves both parameter-inefficient and computationally expensive. To enable efficient adaptation, adapter modules inject lightweight trainable parameters into PLMs while keeping the original weights frozen. These modules employ a bottleneck architecture to minimize trainable parameters, consisting of three layers: 1) a down-projection layer that reduces the original $d$-dimensional representation $x$ to a lower dimension $r$, 2) a nonlinear activation function $f(\cdot)$, and 3) an up-projection layer that restores the features to dimension $d$. The architecture ensures near-zero initialization of the projection layers while maintaining a skip connection to preserve the original features during initial training stages. For an input representation $x\in\mathbb{R}^{d}$, the adapter’s output $y\in\mathbb{R}^{d}$ can be formalized as:
@@
Building upon the standard Adapter’s bottleneck architecture, OA-Adapter introduces two structural modifications: 1) the removal of bias terms in projection layers to create a bias-free parameter space containing only linear transformations, and 2) the replacement of static non-linear activations with a trainable diagonal masking matrix $\Gamma$ that dynamically adjusts the bottleneck dimension, as detailed in Section 2.2. These modifications enable the enforcement of orthogonal parameter subspace constraints, as detailed in Section 2.3, to mitigate cross-task interference and co-optimization of budget adaptation with continual learning in a single training phase. Specifically, the forward computation of OA-Adapter operates as follows:
@@ 2.2 Dynamic Bottleneck Dimension Adaptation
The derivative $\frac{\partial\gamma_{i}}{\partial g_{i}}$ of the soft thresholding function equals $\mathrm{sign}(g_{i})$ when $|g_{i}|>\tau$, but critically becomes 00 when $|g_{i}|\leq\tau$. This implies that deactivated dimensions (where $|g_{i}|\leq\tau$) produce zero gradients in Equation (5) due to $\frac{\partial\gamma_{i}}{\partial g_{i}}=0$, blocking gradient updates through $g_{i}$. With a fixed threshold $\tau$, such dimensions would remain permanently disabled throughout training. Crucially, our method implements $\tau$ as a learnable parameter shared across all dimensions within each OA-Adapter module. Thus, the gradient of $\tau$ with respect to the total loss $\mathcal{L}$ is derived through chain rule as:
@@ 2.3 Orthogonal Parameter Subspace Constraints for Continual Learning
Catastrophic forgetting arises when task-specific adaptations overwrite parameters critical for previous tasks. To mitigate this, we introduce orthogonality constraints that enforce parameter updates across tasks to occupy mutually independent subspaces. Let $\Delta\Phi_{k}=\mathcal{W}_{2}^{(k)}\Gamma^{(k)}\mathcal{W}_{1}^{(k)}$ represent the OA-Adapter’s parameter update for the $k$-th task. We have:
$\Delta\Phi_{k}={\widetilde{\mathcal{W}}_{2}^{(k)}}\cdot\mathcal{W}_{1}^{(k)}=(\mathcal{W}_{2}^{(k)}\cdot{\Gamma^{(k)}})\cdot\mathcal{W}_{1}^{(k)}$ (8)
Here, the columns of $\widetilde{\mathcal{W}}_{2}^{(k)}$ serve as orthogonal basis vectors spanning the parameter update subspace for the $k$-th task, while $\mathcal{W}_{1}^{(k)}$ determines how these basis vectors are combined. We therefore formally define the task-specific parameter subspace as the column space of $\widetilde{\mathcal{W}}_{2}^{(k)}$, which intrinsically aligns with the activated dimensions for the $k$-th task through the dimension-selective masking operation of $\Gamma^{(k)}$. Thus, we enforce strict orthogonality to new OA-Adapter parameters across sequential tasks, ensuring new task adaptations occupy parameter subspaces orthogonal to previous tasks’ frozen parameter subspaces. Formally, the constraints for the $t$-th task is defined as:
@@
The columns of $\widetilde{\mathcal{W}}_{2}^{(t)}$ inherit directional properties from $\mathcal{W}_{2}^{(t)}$, ensuring orthogonal relationships persist regardless of dynamic dimension activation patterns. These asymmetric orthogonality constraints enable simultaneous optimization of dynamic bottleneck dimension adaptation and historical knowledge preservation. To formalize this approach, we incorporate an orthogonality regularization term into the optimization objective. Specifically, the pairwise orthogonality loss between current task $t$ and each historical task $s<t$ is quantified as:
@@
Minimizing the loss term $\mathcal{L}_{\mathrm{orth}}^{(s,t)}$ drives the inner product $\langle{\mathcal{W}}_{2}^{(t)}[:,i],\widetilde{\mathcal{W}}_{2}^{(s)}[:,j]\rangle$ toward zero, enforcing parameter subspace orthogonality. The complete training objective, integrating both task-specific performance and orthogonality constraints, is formulated as:
@@
where $\mathcal{L}_{\mathrm{task}}^{(t)}$ represents the primary loss for task $t$, and $\lambda_{\mathrm{orth}}$ is a hyperparameter controlling the strength of orthogonal regularization.
3 Experiments

### v2

# Adaptive Budget Allocation for Orthogonal-Subspace Adapter Tuning in LLMs Continual Learning

@@ Abstract
Large language models (LLMs) often suffer from catastrophic forgetting in continual learning (CL) scenarios, where performance on previously learned tasks degrades severely while training on sequentially arriving tasks. Although pioneering CL approaches using orthogonal subspaces can mitigate task interference, they typically employ fixed budget allocation, neglecting the varying complexity across tasks and layers. Besides, recent budget-adaptive tuning methods for LLMs often adopt multi-stage paradigms that decouple optimization and budget allocation. Such decoupling results in potential misalignment, which hinders those approaches’ practical application in CL scenarios. To address these limitations, we propose OA-Adapter, a novel parameter-efficient approach for continual learning in LLMs that unifies dynamic budget adaptation with orthogonal subspace learning in an end-to-end training stage. Specifically, OA-Adapter introduces a dynamic bottleneck dimension adaptation mechanism that simultaneously allocates an efficient parameter budget and optimizes task objectives without misalignment. To effectively preserve previously acquired knowledge while coordinating with the dynamic budget allocation, orthogonal constraints are applied specifically between the parameter subspace of the current task and the dynamically allocated parameter subspaces of historical tasks. Experimental results on continual learning benchmarks demonstrate that OA-Adapter outperforms state-of-the-art methods in both accuracy and parameter efficiency. OA-Adapter achieves higher average accuracy while using $58.5\%$ fewer parameters on the standard CL benchmark, and maintains its advantages on two larger benchmarks comprising 15 tasks.
@@ 1 Introduction
Recent advances in large language models (LLMs) have transformed artificial intelligence by demonstrating remarkable capabilities across diverse domains (Zhou et al., 2024a; Touvron et al., 2023). However, real‐world deployment demands that LLMs continually adapt to evolving user needs and emerging tasks while retaining previously acquired knowledge—a prerequisite for sustainable lifelong learning (Xi et al., 2025; Wang et al., 2023b).
@@
Parameter-efficient fine-tuning (PEFT) methods, such as adapter modules (Houlsby et al., 2019) and low-rank adaptation (LoRA) (Hu et al., 2022), enable task‐specific adaptation by updating only $0.01\%-4\%$ of model weights (Balne et al., 2024)). While originally designed to reduce computational costs for single‐task tuning (Zhou et al., 2024b; Chen et al., 2024), these methods struggle in continual learning with sequentially arriving tasks. Sequentially tuning to arriving tasks induces catastrophic forgetting‐severe degradation of performance on prior tasks (McCloskey & Cohen, 1989). One intuitive solution is to store task-specific adapters for each new task. However, this approach consumes substantial storage resources and leads to considerable inflexibility in multi-task deployments. Alternatively, retraining models with archived historical and additional new data necessitates frequent model updates and large data repositories (Shi et al., 2024). Both strategies are prohibitively costly and impractical in resource-constrained environments.
@@
To efficiently adapt LLMs to downstream tasks while preserving previously acquired knowledge, researchers proposed continual learning (CL) methods. However, most CL approaches operate within shared parameter spaces across tasks, which inherently induces cross-task interference (Wang et al., 2024; Shi et al., 2024; Chaudhry et al., 2019; Shi & Wang, 2023; Rebuffi et al., 2017; Aljundi et al., 2018; Schwarz et al., 2018; Rongali et al., 2020; Lin et al., 2022). Furthermore, unlike conventional CL (e.g., incremental image classification) that typically handle tasks with limited distribution shifts , CL of LLMs often deals with substantially divergent task distributions, thus significantly amplifying interference when tuning in a shared parameter space. Some architecture-based methods further construct task-specific parameters to resolve this problem, but they heavily rely on explicit task ID during inference (Razdaibiedina et al., 2023; Jang et al., 2022; Jin et al., 2022; Li & Lee, 2024; Yan et al., 2023). Recent researches in orthogonal subspace learning (Farajtabar et al., 2020; Guo et al., 2022; Wang et al., 2023a) offer a promising alternative by restricting task-specific updates to mutually orthogonal parameter subspaces to eliminate interference, without task ID dependency during inference. However, existing methods typically rely on a fixed budget allocation, assigning the same subspace dimensionality to different task and layer. This rigid strategy overlooks the heterogeneity of task complexity and layer-specific adaptation needs, leading to inefficient parameter utilization, allocating excessive resource to simple tasks while under-allocating resources to more complex ones. Such inflexible allocation hinder LLMs’ continuous adaptation capabilities in practice.
@@
To achieve dynamic budget allocation, emerging budget-adaptive PEFT methods like AdaLoRA (Zhang et al., 2023), ElaLoRA (Chang et al., 2025), and DiffoRA (Jiang et al., 2025) proposed multi-stage paradigms with sequential optimization and budget adjustment phases. Such decoupled optimization may create misalignment between fine-tuning objectives and budget allocation. Moreover, the inherent complexity of multi-stage designs introduces substantial computational overhead and engineering challenges, limiting their practicality for continual learning systems.
@@
We propose OA-Adapter to address these issues, a novel parameter-efficient approach for continual learning in LLMs. Instead of manually assigning a fixed budget, OA-Adapter automatically adjusts the parameter budget for each task and layer based on the task difficulty and model capacity.
@@
  • We propose OA-Adapter, a novel parameter-efficient approach for continual learning in LLMs that unifies dynamic budget adaptation with orthogonal subspace learning in an end-to-end training stage. To the best of our knowledge, this is the first work to integrate budget adaptation into parameter-efficient fine-tuning for continual learning in LLMs.
@@
  • We demonstrate that the parameter budget requirements in continual learning vary with task difficulty and position in the training sequence, and that the budget requirements of parameters in different layers vary, which emphasizes the necessity of our approach.
  • Experimental results demonstrate that OA-Adapter outperforms state-of-the-art methods in both accuracy and parameter efficiency. OA-Adapter achieves higher average accuracy with $58.5\%$ fewer parameters on the standard CL benchmark and maintains its advantages on two larger benchmarks comprising 15 tasks.
2 Related Work.
Continual Learning for LLMs.
Existing continual learning (CL) techniques for LLMs are mainly based on replay, regularization, and architecture modification, which have been most extensively applied in this context (Wang et al., 2024; Shi et al., 2024). Replay-based methods (Chaudhry et al., 2019; Shi & Wang, 2023; Rebuffi et al., 2017) store a subset of past data to rehearse on previous tasks during new training. However, this approach raises significant privacy issues and imposes high storage and computational costs, especially for LLMs. Regularization-based methods (Aljundi et al., 2018; Schwarz et al., 2018; Rongali et al., 2020; Lin et al., 2022) introduce a regularization term that penalizes significant weight deviations across tasks. However, their computational complexity grows rapidly with the number of tasks, and performance on previous tasks deteriorates significantly. These above approaches primarily focus on learning all incremental tasks within a shared parameter space, which is a major contributor to task interference. In contrast, many architecture-based methods (Razdaibiedina et al., 2023; Jang et al., 2022; Jin et al., 2022; Li & Lee, 2024; Yan et al., 2023) address this challenge by incorporating task-specific components, isolated parameters, or dedicated pathways within the model. They essentially learn task-specific expert modules for different tasks, so rely heavily on explicit task ID during inference. Recent advancements have explored a promising direction that can retain generalization capacity while reducing task interference without requiring explicit task ID. These methods (Farajtabar et al., 2020; Guo et al., 2022; Wang et al., 2023a) constrain weight updates for each task to lie within mutually orthogonal subspaces of the high-dimensional parameter space. By doing so, they effectively decouple the optimization directions of different tasks, thereby mitigating task interference. However, a common limitation of existing orthogonal subspace methods is the fixed parameter budget for all tasks and layers.
Adaptive Budget Tuning.
Existing budget-adaptive tuning methods for LLMs commonly adopt a multi-stage design where budget adaptation is a separate stage from optimization. For example, DiffoRA (Jiang et al., 2025) trains a LoRA module for each layer in the first stage, and then selects a subset of these modules based on the ”Difference-aware Adaptive Matrix” in a subsequent stage. ElaLoRA (Chang et al., 2025) involves a ”Warm-up” fine-tuning stage, followed by a ”Dynamic Rank Adjustment” stage to adapt LoRA module ranks, and finally a ”Stabilization” fine-tuning stage. AdaLoRA (Zhang et al., 2023) injects rank pruning operations after iterations by applying SVD to selectively retain low-rank components. While these methods achieve strong performance on single-task datasets, their multi-stage design introduces substantial computational and engineering overhead that compromises practicality and scalability, while creating misalignment between optimization objectives and budget adaptation criteria that ultimately hinders the achievement of truly optimal solutions. And their design is restricted to unidirectional rank reduction, which does not allow for bidirectional budget adaptation. These limitations pose significant challenges when extending such methods to continual learning, which demands efficient, unified, and adaptive training processes for sequential tasks. Moreover, although recent methods (Liu et al., 2024; Chang et al., 2025; Ding et al., 2023) provide strong interpretability, they rely on task-specific heuristic configurations rather than adapting to the inherent characteristics of sequentially arriving tasks.
@@ 3 Methodology
In this section, we introduce OA-Adapter, a novel framework for continual learning in LLMs that simultaneously improves parameter efficiency and mitigates catastrophic forgetting in an end-to-end training stage, as illustrated in Figure 1. We first describe its architectural design, including core components and computation flow. Then, we analyze the mathematical foundations of the dynamic bottleneck dimension adaptation, demonstrating how trainable thresholds enable bidirectional activation and deactivation of dimensions during an end-to-end training phase. Finally, we formalize the orthogonal parameter subspace constraints mechanism and explain how it works in concert with the dynamic bottleneck dimension adaptation to achieve parameter-efficient continual learning.
Figure 1: The OA-Adapter framework for LLM continual learning. Each task-specific OA-Adapter module (task $t$) comprises three core components: (1) a down-projection layer $\mathcal{W}_{1}^{(t)}$, (2) a diagonal mask $\Gamma^{(t)}$ with trainable threshold $\tau^{(t)}$, (3) and an up-projection layer $\mathcal{W}_{2}^{(t)}$. The dynamic masking mechanism enables bidirectional dimension adaptation through activation/deactivation of latent dimensions. Orthogonal subspace constraints are enforced between the column space of the $t$-th task parameters $\mathrm{Col}(\mathcal{W}_{2}^{(t)})$ and the dynamically allocated parameter subspaces of historical tasks $\mathrm{Col}(\widetilde{\mathcal{W}}_{2}^{(s)})$ (for s<t\ s<t). Here, $\widetilde{\mathcal{W}}_{2}^{(s)}$ incorporates only the activated dimensions from the $s$-th task.
@@ 3.1 Module Structure
To enable efficient adaptation of pre-trained language models (PLMs) to downstream tasks, adapter modules inject lightweight trainable parameters into PLMs while keeping the original weights frozen. These modules employ a bottleneck architecture to minimize trainable parameters, consisting of a down-projection layer that reduces the original $d$-dimensional representation $x$ to a lower dimension $r$, a nonlinear activation function $f(\cdot)$, and an up-projection layer that restores the features to dimension $d$. The architecture ensures near-zero initialization of the projection layers while maintaining a skip connection to preserve the original features during initial training stages. For an input representation $x\in\mathbb{R}^{d}$, the output $y\in\mathbb{R}^{d}$ can be formalized as:
@@
Building upon the standard Adapter’s bottleneck architecture, OA-Adapter introduces two structural modifications: 1) the removal of bias terms in projection layers to create a bias-free parameter space containing only linear transformations, and 2) the replacement of static non-linear activations with a trainable diagonal masking matrix $\Gamma$ that dynamically adjusts the bottleneck dimension, as detailed in Section 3.2. These modifications enable the enforcement of orthogonal parameter subspace constraints, as detailed in Section 3.3, to mitigate cross-task interference and co-optimization of budget adaptation with continual learning in an end-to-end training phase. Specifically, the forward computation of OA-Adapter operates as follows:
@@ 3.2 Dynamic Bottleneck Dimension Adaptation
The derivative $\frac{\partial\gamma_{i}}{\partial g_{i}}$ of the soft thresholding function equals $\mathrm{sign}(g_{i})$ when $|g_{i}|>\tau$, but critically becomes 0 when $|g_{i}|\leq\tau$. This implies that deactivated dimensions (where $|g_{i}|\leq\tau$) produce zero gradients in Equation (5) due to $\frac{\partial\gamma_{i}}{\partial g_{i}}=0$, blocking gradient updates through $g_{i}$. With a fixed threshold $\tau$, such dimensions would remain permanently disabled throughout training. Crucially, our method implements $\tau$ as a learnable parameter shared across all dimensions within each OA-Adapter module. Thus, the gradient of $\tau$ with respect to the total loss $\mathcal{L}$ is:
@@ 3.3 Orthogonal Parameter Subspace Constraints for Continual Learning
Catastrophic forgetting arises when subsequent adaptations overwrite parameters critical for previous tasks. To mitigate this, we introduce orthogonality constraints that enforce parameter updates across tasks to occupy mutually independent subspaces. Let $\Delta\Phi_{k}$ represent the OA-Adapter’s parameter for the $k$-th task. We have:
$\Delta\Phi_{k}=\mathcal{W}_{2}^{(k)}\cdot\Gamma^{(k)}\cdot\mathcal{W}_{1}^{(k)}={\widetilde{\mathcal{W}}_{2}^{(k)}}\cdot\mathcal{W}_{1}^{(k)}$ (8)
Here, the columns of $\widetilde{\mathcal{W}}_{2}^{(k)}$ serve as orthogonal basis vectors spanning the parameter update subspace for the $k$-th task, while $\mathcal{W}_{1}^{(k)}$ determines how these basis vectors are combined. We therefore formally define the task-specific parameter subspace as the column space of $\widetilde{\mathcal{W}}_{2}^{(k)}$, which intrinsically aligns with the activated dimensions for the $k$-th task through the dimension-selective masking operation of $\Gamma^{(k)}$. Thus, we enforce strict orthogonality to new OA-Adapter parameters across sequential tasks, ensuring new task adaptations occupy parameter subspaces orthogonal to previous tasks’ frozen parameter subspaces. Formally, the constraints for the $t$-th task are defined as:
@@
The columns of $\widetilde{\mathcal{W}}_{2}^{(t)}$ inherit directional properties from $\mathcal{W}_{2}^{(t)}$, ensuring orthogonal relationships persist regardless of dynamic dimension activation patterns. These asymmetric orthogonality constraints enable simultaneous optimization of dynamic bottleneck dimension adaptation and historical knowledge preservation. We detail the analysis in Appendix A.2. To formalize this approach, we incorporate an orthogonality regularization term into the optimization objective. Specifically, the pairwise orthogonality loss between current task $t$ and each historical task $s<t$ is quantified as:
@@
Minimizing the loss term $\mathcal{L}_{\mathrm{orth}}^{(s,t)}$ drives the inner product $\langle{\mathcal{W}}_{2}^{(t)}[:,i],\widetilde{\mathcal{W}}_{2}^{(s)}[:,j]\rangle$ toward zero, enforcing parameter subspace orthogonality. The complete training objective, integrating both task-specific performance term $\mathcal{L}_{\mathrm{task}}^{(t)}$ and orthogonality constraints term $\mathcal{L}_{\mathrm{orth}}^{(s,t)}$, is formulated as:
@@
where $\lambda_{\mathrm{orth}}$ is a hyperparameter controlling the strength of orthogonal regularization.
4 Experiments and Analysis